# Can Implicit Bias Imply Adversarial Robustness?

### v1

@@ 3 Main Results on Adversarial Robustness
Comparison with Frei et al. (2023). Frei et al. (2023) considers a similar setting as ours with only a slight difference: their data distribution differs from ours by a scaling factor of $\sqrt{D}$ on the input data (they also allow tiny correlations between two subclass centers) and they considers shallow ReLU network with bias. They show that any trained network by gradient flow is non-robust against $\mathcal{O}\left(\frac{1}{\sqrt{K}}\right)$ attack, which covers any initialization, but their analysis does not characterize what the classifier the trained network corresponds to. While we consider specifically small initialization, we can at least conjecture, and numerically verify, what the corresponding classifier is at the end of the training. In addition, Frei et al. (2023) show the existence of $\mathcal{O}(1)$-robust ReLU network if neurons are aligned with subclasses and there is some carefully chosen bias. While it already sheds light on the need for learning subclasses, their proposed network can not be found by gradient flow. On the contrary, our proposed robust $F^{(p)}(\boldsymbol{x})$ can be obtained by gradient flow.
@@ 4 Discussion on Gradient Flow Training
In the previous section, we conjectured that under gradient flow starting from a small initialization, a vanilla ReLU network favors learning average directions $\bar{\boldsymbol{\mu}}_{+}$ and $\bar{\boldsymbol{\mu}}_{-}$ of subclass centers, while a pReLU network favors learning every subclass centers $\boldsymbol{\mu}_{k},\ k\in[K]$, resulting in a significant difference between these two networks in terms of adversarial robustness. In this section, we provide a theoretical explanation of such alignment preferences in the scope of implicit bias of gradient flow training with small initialization (Maennel et al., 2018; Boursier & Flammarion, 2024). We remind the reader that gradient flow training is described in Section 2.
Figure 1: Visualizing the training of a pReLU network under small initialization. The dataset has its positive class sampled from two subclasses. (a) At initialization, all neurons have small norms and point toward random directions; When $p=1$, (b) During the alignment phase, the neuron directions are aligned with either of the average class centers $\bar{\boldsymbol{\mu}}_{+}$ and $\bar{\boldsymbol{\mu}}_{-}$; (c) During the second phase, neurons keep the alignment with $\bar{\boldsymbol{\mu}}_{+}$ and $\bar{\boldsymbol{\mu}}_{-}$ while growing their norms; When $p=3$, (d) neurons learn subclass centers during alignment phase and (e) keep the alignment in the second phase. Note: the neurons pointing toward directions other than class/subclass centers are those not activated by any of the data points and have small norms throughout training.
@@ 4.2 Preliminary theoretical analysis on neuron alignment in pReLU networks
$\;\max_{j}\left\lVert\frac{d}{dt}\frac{\boldsymbol{w}_{j}(t)}{\|\boldsymbol{w}_{j}(t)\|}-\mathrm{sign}(v_{j}(0))\mathcal{P}^{\perp}_{\boldsymbol{w}_{j}(t)}\boldsymbol{x}^{(p)}(\boldsymbol{w}_{j}(t))\right\rVert$
$\;\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad=\mathcal{O}\left(\epsilon k\sqrt{h}\right)\,,$

### v2

@@ 3 Main Results on Adversarial Robustness
Comparison with prior work. Frei et al. (2023) consider a similar setting as ours with three minor differences. Firstly, their data distribution differs from ours by a scaling factor of $\sqrt{D}$ on the input data. Second, they allow for tiny correlations between the two subclass centers. Third, they consider shallow ReLU networks with bias. They show that any network trained by gradient flow is non-robust against $\mathcal{O}\left(\frac{1}{\sqrt{K}}\right)$ attacks, which covers any initialization, but their analysis does not characterize what classifier the trained network corresponds to. While we consider specifically small initialization, we can at least conjecture, and numerically verify, what the corresponding classifier is at the end of the training. In addition, Frei et al. (2023) show the existence of $\mathcal{O}(1)$-robust ReLU network if neurons are aligned with subclasses and there is some carefully chosen bias. While it already sheds light on the need for learning subclasses, their proposed network can not be found by gradient flow. On the contrary, our proposed robust $F^{(p)}$ can be obtained by gradient flow.
@@ 4 Discussion on Gradient Flow Training
In Section 3, we conjectured that under gradient flow starting from a small initialization, a vanilla ReLU network favors learning average directions $\bar{\boldsymbol{\mu}}_{+}$ and $\bar{\boldsymbol{\mu}}_{-}$ of subclass centers, while a pReLU network favors learning every subclass centers $\boldsymbol{\mu}_{k},\ k\in[K]$, resulting in a significant difference between these two networks in terms of adversarial robustness. In this section, we provide a theoretical explanation of such alignment preferences in the scope of implicit bias of gradient flow training with small initialization (Maennel et al., 2018; Boursier & Flammarion, 2024). We remind the reader that gradient flow training is described in Section 2.
@@ 4.2 Preliminary theoretical analysis on neuron alignment in pReLU networks
$\;\max_{j}\left\lVert\frac{d}{dt}\frac{\boldsymbol{w}_{j}(t)}{\|\boldsymbol{w}_{j}(t)\|}-\mathrm{sign}(v_{j}(0))\mathcal{P}^{\perp}_{\boldsymbol{w}_{j}(t)}\boldsymbol{x}^{(p)}(\boldsymbol{w}_{j}(t))\right\rVert=\mathcal{O}\left(\epsilon k\sqrt{h}\right)\,,$